# Temporal Cooperative Games

### v3

@@ 5 Solutions Satisfying Shapley-inspired Properties
In this section, we introduce two solution concepts, one each in $\Phi$ and $\Psi$ respectively. We show that both these solution concepts uniquely satisfy certain sets of desired properties defined in the previous section.
@@
$\textsc{Ext-Shap}_{i}=\frac{1}{n!}\sum_{\pi\in\Pi:|\pi|=n}\textsc{MargSol}_{i}(\pi)=\frac{1}{n!}\sum_{\pi\in\Pi:|\pi|=n}(v(\pi_{i}+i)-v(\pi_{i})).$ (11)
@@
Proof.
Part (a): since $\phi$ satisfies SE, we have $\sum_{i\in P(\pi)}\phi_{i}(\pi,v)=v(\pi),\forall\pi\in\Pi$. Averaging this over all full length sequences, we get $\forall\pi\in\Pi$
$=\sum_{i\in N}\overline{\phi}_{i}(\pi,v)$
The first equality comes via changing the order of the summation and the second equality is from the definition of the reduced solution concept. Hence, $\overline{\phi}$ satisfies EE.
$\frac{1}{n!}\sum_{\pi\in\Pi:P(\pi)=N}(\phi_{i}(\pi,u)+\phi_{i}(\pi,w))$ $=\frac{1}{n!}\sum_{\pi\in\Pi:P(\pi)=N}\phi_{i}(\pi,u+w),\>\forall\pi\in\Pi,i\in N$
Hence, $\overline{\phi}$ satisfies ENP. ∎
@@
Proof.
$\sum_{\pi\in\Pi}\alpha_{\pi}u_{\pi}(\pi_{0})=\sum_{\pi\in\Pi,\pi\sqsubset\pi_{0}}\alpha_{\pi}u_{\pi}(\pi_{0})+\alpha_{\pi_{0}}u_{\pi_{0}}(\pi_{0})+\sum_{\pi\in\Pi,\pi\not\sqsubset\pi_{0}}\alpha_{\pi}u_{\pi}(\pi_{0})=0.$
[…]
@@ 5.1 Uniqueness of MargSol
Proof.
$=(v(\pi(1))-v(\emptyset))+(v(\pi(1)\pi(2))-v(\pi(1)))+(v(\pi(1)\pi(2)\pi(3))-v(\pi(1)\pi(2)))+$
$\ldots+(v(\pi(1)\pi(2)\pi(3)\cdots\pi(|P(\pi)|))-v(\pi(1)\pi(2)\cdots\pi(|P(\pi)|-1)))=v(\pi).$
Where $\pi(k)$ is the agent at the $k^{\textup{th}}$ position of $\pi$.
For each carrier game, we know the structure of the unique solution concept that satisfies SE and SNP from Lemma˜4. Note that MargSol satisfies all the three properties for every game $\langle N,v\rangle$, hence it must satisfy them for the carrier games as well. Suppose there exists a different solution concept $\phi$ that also satisfies the same three properties. We will prove that $\phi$ must be MargSol.
From Lemma˜4 and the discussion above, we conclude that $\phi$ must be same as MargSol for the carrier game $\langle N,u_{\pi,\alpha}\rangle$. Hence we have
$\textsc{MargSol}_{i}(\pi^{\prime},u_{\pi,\alpha_{\pi}})$ $=\phi_{i}(\pi^{\prime},u_{\pi,\alpha_{\pi}}),\ \forall i\in N,\forall\pi,\pi^{\prime}\in\Pi,$
$\Rightarrow$ $\sum_{\pi\in\Pi}\textsc{MargSol}_{i}(\pi^{\prime},u_{\pi,\alpha_{\pi}})$ $=\sum_{\pi\in\Pi}\phi_{i}(\pi^{\prime},u_{\pi,\alpha_{\pi}}),\ \forall i\in N,\forall\pi^{\prime}\in\Pi,$
$\Rightarrow$ $\textsc{MargSol}_{i}\left(\pi^{\prime},\sum_{\pi\in\Pi}u_{\pi,\alpha_{\pi}}\right)$ $=\phi_{i}\left(\pi^{\prime},\sum_{\pi\in\Pi}u_{\pi,\alpha_{\pi}}\right),\ \forall i\in N,\forall\pi^{\prime}\in\Pi,$
$\Rightarrow$ $\textsc{MargSol}_{i}\left(\pi^{\prime},v\right)$ $=\phi_{i}\left(\pi^{\prime},v\right),\ \forall i\in N,\forall\pi^{\prime}\in\Pi.$
The first implication holds by summing over all $\pi\in\Pi$. The second implication is by SA for both solution concepts. The third implication is by Equation˜12. Since $v$ was arbitrary, we conclude that $\phi$ is the same as MargSol. ∎
@@ 5.2 Uniqueness of Ext-Shap
Ext-Shap is an averaged version of MargSol (Equation˜11) and hence the mapping to Ext-Shap may not be unique. Hence, even though MargSol uniquely satisfies SE, SA, and SNP (Theorem˜2), that does not automatically imply that Ext-Shap will uniquely satisfy EE, EA, and ENP. In this section, we show that indeed this implication is true via a proof similar yet independent of that of Theorem˜2.
Lemma 5.
Proof.
Theorem 3.
@@
Proof.
($\Rightarrow$) This direction of the proof follows in similar lines of that of Theorem˜2. Assume that there exists an extended solution concept $\psi\neq\textsc{Ext-Shap}{}$ that also satisfies the three given properties.
For each carrier game, we know the structure of the unique solution concept that satisfies SE and SNP from Lemma˜5. Note that Ext-Shap satisfies all the three properties for every game $\langle N,v\rangle$, hence it must satisfy them for the carrier games as well. Hence $\psi$ must be same as Ext-Shap for the carrier game $\langle N,u_{\pi,\alpha}\rangle$. Hence we have
The first implication holds by summing over all $\pi\in\Pi$. The second implication is by EA for both solution concepts. The third implication is by Equation˜12. Since $v$ was arbitrary, we conclude that $\psi$ is the same as Ext-Shap. ∎
Remark.
The Shapley value equivalents in the TCG world, MargSol and Ext-Shap, satisfy ‘almost’ similar uniqueness properties like the Shapley value. However, a careful reader can spot the difference since Shapley value was unique for four (and not three as we did here) properties which also included symmetry. The TCG-equivalent definitions of symmetry are given in Definitions˜11 and 12. In the following section, we show why these two properties were not necessary to obtain the uniqueness theorems Theorems˜2 and 3.
@@ 6 Why Symmetry Is Not Necessary?
Properties in the TCG framework are more restrictive as they make claims for all sequences, making them stronger than claims for all sets, as in the classical cooperative games. Hence, three properties are sufficient to ensure unique solution concepts like MargSol and Ext-Shap. In the following, we show that MargSol and Ext-Shap satisfy SS (Definition˜11) and ES (Definition˜12) respectively defined in the spirit of TCGs. These result show that the three other properties imply the symmetry property.
@@
Remarks.
In classical cooperative games, the Shapley value is the only solution concept satisfying efficiency, additivity, null player property and symmetry. Symmetry is required to make the solution concept unique. The difference in our setting lies in the definition of the null player property for TCGs.
In the classical setting, a carrier game is defined with respect to a critical set of players. When the set is present, only then a coalition generates unit positive worth. Dividing the worth among the non-null players requires a property like symmetry.
In TCG framework, a carrier game corresponds to a critical sequence of players. When this critical sequence happens to be a prefix of a given sequence, only then it can generate unit positive worth. As a result, only the last player of the critical sequence is non-null, so symmetry is not required. Only one player gets all the worth.
@@ 7 SeqShare and Shapley-compliant: Best of Both Worlds
SeqShare uniquely satisfies three most desirable properties for TCGs: OIR, SE, and I4OA. On the other hand, Ext-Shap uniquely characterizes the three most desirable Shapley-equivalent properties: EA, EE, and ENP. A solution concept that satisfies the temporal properties of TCGs and simultaneously has its reduced version satisfying the Shapley properties would therefore be the ideal candidate in a TCG. In this section we show that such a “best of both worlds” solution concept does not exist, even in special games like convex or simple games (Maschler et al., 2020, Chapter 19).
@@
Proof.
$\phi_{i}(\pi)\leqslant\phi_{i}(\pi^{*}(v)),\ \forall\pi\in\Pi$
Averaging over all possible $\pi\in\Pi$ such that $P(\pi)=N$, we get
$\overline{\phi}_{i}=\frac{\textstyle 1}{\textstyle n!}\sum_{\pi\in\Pi:P(\pi)=N}\phi_{i}(\pi)<\frac{\textstyle 1}{\textstyle n!}\sum_{\pi\in\Pi:P(\pi)=N}v(\pi_{i}+i)-v(\pi_{i})=\textsc{Ext-Shap}_{i}.$ (14)
Hence, we conclude that Equation˜13 is necessary to ensure that a solution from SeqShare reduces to Ext-Shap. In the example given below we show a game that does not satisfy Equation˜13 and therefore, there does not exist any solution concept $\phi\in\textsc{SeqShare}{}$ that has a reduction to Ext-Shap.
@@
Consider the TCG shown in Figure˜4 with $a=7,b=3$. Clearly, $v(\pi*)=8$. The basis solution (Equation˜1) dictates that $x_{1}\geqslant v(1)=1,x_{2}\geqslant v(2)=3,x_{3}\geqslant v(3)=4$ and $x_{1}+x_{2}+x_{3}=8$, which implies that the basis solution must be $x_{1}=1,x_{2}=3,x_{3}=4$. This is the only possible basis solution and therefore any $\phi\in\textsc{SeqShare}{}$ assigns $\phi_{i}(\pi^{*})=x_{i}$. Routine calculations for Player 1 gives $\textsc{Ext-Shap}{}_{1}=\frac{8}{6}$. However, $\phi_{1}(\pi^{*})=1<\frac{8}{6}=\textsc{Ext-Shap}{}_{1}$, which violates Equation˜13. ∎
Remark. Since no solution in SeqShare can be reduced to Ext-Shap, no solution in SeqShare can be MargSol. In particular, MargSol satisfies SE by construction. Since we are considering only monotone games, the marginal contribution, $v(\pi_{i}+i)-v(\pi_{i})$, is always non-negative for every player $i\in N$, and remains so even after the arrival of future agents, which ensures OIR. Hence, the only property that MargSol violates is I4OA. The following example illustrates this. This rules out the possibility of obtaining a solution concept that satisfies both set of properties in both the original and extended spaces.
Example 4.
Consider the TCG given in Figure˜5
12v(1)=3v(1)=3v(12)=8v(12)=8
21v(2)=1v(2)=1v(21)=5v(21)=5
Figure 5: MargSol violates I4OA
In this example $\pi^{*}=12$. Consider the MargSol solution:
$\textsc{MargSol}_{1}(1)=1,\textsc{MargSol}{}_{2}(1)=0,\textsc{MargSol}{}_{2}(2)=3,\textsc{MargSol}{}_{1}(2)=0,$
$\textsc{MargSol}{}_{1}(21)=4,\textsc{MargSol}{}_{2}(21)=1,\textsc{MargSol}{}_{1}(12)=3,\textsc{MargSol}{}_{2}(12)=5.$
[…]
@@
Figure 6: Impossibilities of Shapley inspired properties and solutions in SeqSh.
@@ 7.1 Convex Games
In classical cooperative games, Shapley value exhibit some nice properties in the class of convex games. One such property is that for convex games, the Shapley Value always lies in the core. Since the basis conditions (Equation˜1) are analogous to the core conditions in classical games, we check for a similar result in the case of TCGs. We also check if there is a solution in SeqShare that reduces to Ext-Shap in this class of games. For that, first we need to define convex TCGs.
@@
A TCG $\langle N,v\rangle$ is convex if the following holds
$v(\pi+i)-v(\pi)\leqslant v(\pi^{\prime}+i)-v(\pi^{\prime}),\ \forall\pi^{\prime},\pi\text{ s.t. }i\notin P(\pi^{\prime})\text{ and }P(\pi)\subseteq P(\pi^{\prime}).$ (15)
@@ A TCG $\langle N,v\rangle$ is convex if the following holds
Consider the convex TCG as shown in Figure˜4 with $a=7,b=4$. The basis solution (Equation˜1) must satisfy $x_{1}\geqslant 1,x_{2}\geqslant 4,x_{3}\geqslant 4$ and $x_{1}+x_{2}+x_{3}=8$, that are clearly impossible. Hence, the given convex TCG has no basis solution. ∎
Our next example shows that even for a convex TCG having a basis solution, there may be no solution in SeqSh that reduces to Ext-Shap.
@@
Consider the TCG as used in ˜3, the game is convex and violates Equation˜13. ∎
In the next section we define simple TCGs and explore the connections between Ext-Shap and SeqShare.
@@ 7.2 Simple Games
In classical cooperative games simple games or ‘0-1’ games are defined as the games where the worth of a coalition can either be 0 or $1$. In TCGs, we follow an analogous definition where the worth of any sequence can either take 0 or $1$ values. We define simple TCGs as follows.
@@
The following example shows that even for simple games there may be no solution in SeqShare that reduces to Ext-Shap.
@@
Consider the simple TCG given in Figure˜7. The unique basis solution (Equation˜1) is $x_{1}=1,x_{2}=0,x_{3}=0$. Therefore any $\phi\in\textsc{SeqShare}{}$ assigns $\phi_{i}(\pi^{*})=x_{i}$. Routine calculations for Player 2 gives $\textsc{Ext-Shap}{}_{2}=\frac{1}{6}$. However, $\phi_{2}(\pi^{*})=0<\frac{1}{6}=\textsc{Ext-Shap}{}_{2}$, which violates Equation˜13. ∎
@@ 8 Conclusion
In this work, we considered the framework of temporal cooperative games (TCGs), where the worth of a coalition depends on the order of agent arrivals. This setting departs from classical cooperative games and requires rethinking fundamental axioms. We proposed three natural temporal properties, namely incentive for optimal arrival (I4OA), online individual rationality (OIR), and sequential efficiency (SE), and showed that together they uniquely characterize a class of reward-sharing mechanisms. We further developed two Shapley-like analogs, one in the sequential world and one in the extended world, and established that they are uniquely characterized by efficiency, additivity, and the null player property. Importantly, we demonstrated a fundamental incompatibility: the Shapley analogs and the mechanisms satisfying I4OA, OIR, and SE are disjoint, and this conflict persists even in restricted settings such as convex and simple TCGs.
@@
Our results highlight a key tension in extending classical cooperative game theory to temporal environments: desirable temporal properties cannot be reconciled with Shapley-like axioms. As future work, an important direction is to investigate whether a unique solution concept can be obtained by imposing additional properties beyond I4OA, OIR, and SE. Another avenue is to identify special classes of temporal cooperative games where Shapley-like properties and temporal properties such as I4OA can coexist. These questions open new opportunities for developing principled solution concepts for cooperative games in sequential environments.

### v4

@@ 5 Solutions Satisfying Shapley-inspired Properties
In this section, we introduce two solution concepts, in $\Phi$ and $\Psi$ respectively. We show that both these solution concepts uniquely satisfy certain sets of desired properties defined in the previous section.
@@
$\begin{split}\textstyle\textsc{Ext-Shap}_{i}=&\textstyle\frac{1}{n!}\sum_{\pi\in\Pi:|\pi|=n}\textsc{MargSol}_{i}(\pi).\end{split}$ (3)
@@
Proof Sketch.  Averaging the sequential property over all full-length sequences gives the desired extended property. $\square$
@@
Proof Sketch.  The proof follows similarly to the classical Shapley characterisation. The set of all carrier games forms a linearly independent set of size $|\Pi|$, thus forming a spanning set for the vector space of all TCGs. $\square$
@@ 5.1 Uniqueness of MargSol
Proof Sketch.  ($\Leftarrow$) If player $i$ is a sequential null player, then for all $\pi,\ \textsc{MargSol}_{i}(\pi)=0$. Additivity (SA) holds by linearity of marginal contributions, and efficiency (SE) follows from telescoping the sum of incremental values across the sequence.
($\Rightarrow$) Any TCG can be written as a sum of carrier games. For each carrier game, MargSol is the unique solution satisfying SE and SNP. If another solution $\phi$ also satisfies SE, SA, and SNP, it must agree with MargSol on all carrier games. By linearity and the basis decomposition, $\phi$ must equal MargSol on the entire game. $\square$
@@ 5.2 Uniqueness of Ext-Shap
Ext-Shap is an averaged version of MargSol (Equation˜3), and therefore the mapping from MargSol to Ext-Shap need not be unique. Consequently, although MargSol uniquely satisfies SE, SA, and SNP (Theorem˜2), this does not automatically imply that Ext-Shap uniquely satisfies EE, EA, and ENP. However, Nowak and Radzik (1994, Theorem 1) established that Ext-Shap indeed uniquely satisfies EE, EA, and ENP. We restate this theorem below and present an alternative proof in the appendix using the framework developed in this paper. Our characterization of order-sensitive solution concepts thus aligns with, and complements, the characterization of order-oblivious solution concepts existing in the literature.
Theorem 3 (Nowak and Radzik (1994)).
@@
Remark. The Shapley value equivalents in the TCG world, MargSol and Ext-Shap, satisfy ‘almost’ similar uniqueness properties like the Shapley value. However, a careful reader can spot the difference since Shapley value was unique for four (and not three as we did here) properties, which also included symmetry. The TCG-equivalent definitions of symmetry are given in Definitions˜11 and 12. In Section˜6, we show why these two properties were not necessary to obtain the uniqueness theorems Theorems˜2 and 3.
@@ 6 Why Symmetry Is Not Necessary?
Properties in the TCG framework are more restrictive as they make claims for all sequences, making them stronger than claims for all sets, as in the classical cooperative games. Hence, three properties are sufficient to ensure unique solution concepts like MargSol and Ext-Shap. In the following, we show that MargSol and Ext-Shap satisfy SS (Definition˜11) and ES (Definition˜12) respectively, defined in the spirit of TCGs. These results show that the three other properties imply the symmetry property.
@@
Remark. In classical cooperative games, the Shapley value is the only solution concept satisfying efficiency, additivity, null player property and symmetry. Symmetry is required to make the solution concept unique. The difference in our setting lies in the definition of the null player property for TCGs.
In the classical setting, a carrier game is defined with respect to a critical set of players. When the set is present, only then does a coalition generate unit positive worth. Dividing the worth among the non-null players requires a property like symmetry.
In TCG framework, a carrier game corresponds to a critical sequence of players. When this critical sequence happens to be a prefix of a given sequence, only then can it generate a unit positive worth. As a result, only the last player of the critical sequence is non-null, so symmetry is not required. Only one player gets all the worth.
@@ 7 SeqShare and Shapley-compliant: Best of Both Worlds
The SeqShare class characterizes three most desirable properties for TCGs: OIR, SE, and I4OA. On the other hand, Ext-Shap uniquely characterizes the three most desirable Shapley-equivalent properties: EA, EE, and ENP. A solution concept that satisfies the temporal properties of TCGs and simultaneously has its reduced version satisfying the Shapley properties would therefore be the ideal candidate in a TCG. In this section, we show that such a “best of both worlds” solution concept does not exist, even in special games like convex or simple games (Maschler et al., 2020, Chapter 19).
@@
Proof Sketch.  Assume $\phi\in\textsc{SeqShare}$ with $\overline{\phi}=\textsc{Ext-Shap}$. Suppose $\phi_{i}(\pi^{*}(v))$ is strictly less than the average marginal contribution. By I4OA, $\phi_{i}(\pi)\leq\phi_{i}(\pi^{*}(v))$ for all $\pi$, so the average $\overline{\phi}_{i}$ is also strictly less than $\textsc{Ext-Shap}_{i}$. This contradicts the assumption $\overline{\phi}=\textsc{Ext-Shap}$, hence no such $\phi$ can exist. $\square$
Hence, we conclude that Equation˜4 is necessary to ensure that a solution from SeqShare reduces to Ext-Shap. In the example given below, we show a game that does not satisfy Equation˜4, and therefore, there does not exist any solution concept $\phi\in\textsc{SeqShare}{}$ that has a reduction to Ext-Shap.
@@
Consider the TCG shown in Figure˜2 with $a=7,b=3$. Clearly, $v(\pi^{*})=8$. The basis solution (Equation˜1) dictates that $x_{1}\geqslant v(1)=1,x_{2}\geqslant v(2)=3,x_{3}\geqslant v(3)=4$ and $x_{1}+x_{2}+x_{3}=8$, which implies that the basis solution must be $x_{1}=1,x_{2}=3,x_{3}=4$. This is the only possible basis solution and therefore any $\phi\in\textsc{SeqShare}{}$ assigns $\phi_{i}(\pi^{*})=x_{i}$. Routine calculations for Player 1 gives $\textsc{Ext-Shap}{}_{1}=\frac{8}{6}$. However, $\phi_{1}(\pi^{*})=1<\frac{8}{6}=\textsc{Ext-Shap}{}_{1}$, which violates Equation˜4. ∎
Remark. Since no solution in SeqShare can be reduced to Ext-Shap, no solution in SeqShare can be MargSol. In particular, MargSol satisfies SE by construction. Since we are considering only monotone games, the marginal contribution, $v(\pi_{i}+i)-v(\pi_{i})$, is always non-negative for every player $i\in N$, and remains so even after the arrival of future agents, which ensures OIR. Hence, the only property that MargSol violates is I4OA. The following example illustrates this. This rules out the possibility of obtaining a solution concept that satisfies both sets of properties in both the original and extended spaces.
@@
Figure 3: Impossibilities of Shapley-inspired properties and solutions in SeqShare(SeqSh).
@@ 7.1 Convex Games
In classical cooperative games, the Shapley value exhibits some nice properties in the class of convex games. One such property is that for convex games, the Shapley Value always lies in the core. Since the basis conditions (Equation˜1) are analogous to the core conditions in classical games, we check for a similar result in the case of TCGs. We also check if there is a solution in SeqShare that reduces to Ext-Shap in this class of games. For that, first we need to define convex TCGs.
@@
A TCG $\langle N,v\rangle$ is convex if the following holds $v(\pi+i)-v(\pi)\leqslant v(\pi^{\prime}+i)-v(\pi^{\prime}),\ \forall\pi^{\prime},\pi\text{ s.t. }i\notin P(\pi^{\prime})\text{ and }P(\pi)\subseteq P(\pi^{\prime}).$
@@
Consider the convex TCG as shown in ˜3 with $a=7,b=4$. The basis solution (Equation˜1) must satisfy $x_{1}\geqslant 1,x_{2}\geqslant 4,x_{3}\geqslant 4$ and $x_{1}+x_{2}+x_{3}=8$, that are clearly impossible. Hence, the given convex TCG has no basis solution. ∎
Our next example shows that even for a convex TCG having a basis solution, there may be no solution in SeqShare that reduces to Ext-Shap.
@@
Consider the TCG as used in ˜3, the game is convex and violates Equation˜4. ∎
Due to space constraints, we define simple TCGs and explore the connections between Ext-Shap and SeqShare in Section˜7.2.
@@ 7.2 Simple Games
In classical cooperative games, simple games or ‘0-1’ games are defined as the games where the worth of a coalition can either be 0 or $1$. In TCGs, we follow an analogous definition where the worth of any sequence can either take 0 or $1$ values. We define simple TCGs as follows.
@@
The following example shows that even for simple games, there may be no solution in SeqShare that reduces to Ext-Shap.
@@
Consider the simple TCG given in Figure˜4. The unique basis solution (Equation˜1) is $x_{1}=1,x_{2}=0,x_{3}=0$. Therefore any $\phi\in\textsc{SeqShare}{}$ assigns $\phi_{i}(\pi^{*})=x_{i}$. Routine calculations for Player 2 gives $\textsc{Ext-Shap}{}_{2}=\frac{1}{6}$. However, $\phi_{2}(\pi^{*})=0<\frac{1}{6}=\textsc{Ext-Shap}{}_{2}$, which violates Equation˜4. ∎
@@ 8 Conclusions and Future Work
TCGs bring in new challenges for the axioms and depart from classical cooperative games. The solution concepts proposed in this paper were order-sensitive. We characterized three natural temporal properties, namely incentive for optimal arrival (I4OA), online individual rationality (OIR), and sequential efficiency (SE), with a class of reward-sharing mechanisms. The Shapley analog for order-sensitive solution concepts uniquely characterizes efficiency, additivity, and the null player property. Importantly, we demonstrated a fundamental incompatibility: the Shapley analogs and the mechanisms satisfying I4OA, OIR, and SE are disjoint, and this conflict persists even in restricted settings such as convex and simple TCGs.
@@
Our results highlight a key tension in extending classical cooperative game theory to generalized characteristic functions: desirable temporal properties cannot be reconciled with Shapley-like axioms. As future work, an important direction is to investigate whether a unique solution concept can be obtained by imposing additional properties beyond I4OA, OIR, and SE. Another avenue is to identify special classes of temporal cooperative games where Shapley-like properties and temporal properties such as I4OA can coexist. These questions open new opportunities for developing principled solution concepts for cooperative games in sequential environments.
Appendix A Appendix to Section˜3
A.1 Proof of Lemma˜1
Proof.
Hence $\phi(\pi^{*})$ satisfies basis conditions. ∎
A.2 proof of Theorem˜1
Proof.
First, observe that for all unit-length sequences, the allocation given by Algorithm˜1 trivially satisfies SE. Property I4OA is satisfied due to the fact that $\phi_{i}(\{i\})=v(\{i\})\leqslant\phi_{i}(\pi^{*}(v)),\forall i\in N$, given by the Basis conditions. OIR is trivial too, since the game is monotone and $\phi_{i}(\emptyset)=0,\forall i\in N$.
Hence, by the induction hypothesis, assume that the forward implication of the theorem is true for all $\pi$ such that $|\pi|=k-1,k\geqslant 2$. We show that the implication is true for every $\pi+i$ of length $k$ as well. Consider the function Improvize. For every $\pi$ of length k1k-1 and the existing partial share $\phi(\pi)$, Improvize allocates the reward to agents in such a way that it is at least $\phi_{i}(\pi)$ (required by OIR) and at most $\phi_{i}(\pi^{*})$ (required by I4OA). This allocation is always possible since the monotonicity of the game ensures $v(\pi+i)\geqslant v(\pi)$ and $v(\pi+i)\leqslant v(\pi^{*})$ (by definition of $\pi^{*}$). Hence, I4OA and OIR are satisfied for each $i\in N$. Also, the construction of the $y$ vectors (reward shares) is such that the sum of rewards of all agents $j\in\pi+i$ equals $v(\pi+i)$, satisfying SE.
Now, set, $\forall i\in P(\pi_{\textup{longshort}})$
$=\phi_{i}(\pi_{\textup{longshort}}).$
Appendix B Appendix to Section˜5
B.1 Proof of Lemma˜2
Proof.
Part (a): since $\phi$ satisfies SE, we have $\sum_{i\in P(\pi)}\phi_{i}(\pi,v)=v(\pi),\forall\pi\in\Pi$. Averaging this over all full-length sequences, we get $\forall\pi\in\Pi$
$=\sum_{i\in N}\overline{\phi}_{i}(\pi,v)$
The first equality comes via changing the order of the summation, and the second equality is from the definition of the reduced solution concept. Hence, $\overline{\phi}$ satisfies EE.
$\frac{1}{n!}\sum_{\pi\in\Pi:P(\pi)=N}(\phi_{i}(\pi,u)+\phi_{i}(\pi,w))$
$=\frac{1}{n!}\sum_{\pi\in\Pi:P(\pi)=N}\phi_{i}(\pi,u+w),\>\forall\pi\in\Pi,i\in N$
Hence, $\overline{\phi}$ satisfies ENP. ∎
B.2 Proof of Lemma˜3
Proof.
$\sum_{\pi\in\Pi}\alpha_{\pi}u_{\pi}(\pi_{0})=$ $\sum_{\pi\in\Pi,\pi\sqsubset\pi_{0}}\alpha_{\pi}u_{\pi}(\pi_{0})+\alpha_{\pi_{0}}u_{\pi_{0}}(\pi_{0})$
$+\sum_{\pi\in\Pi,\pi\not\sqsubset\pi_{0}}\alpha_{\pi}u_{\pi}(\pi_{0})=0.$
[…]
B.3 Proof of Theorem˜2
Proof.
$=(v(\pi(1))-v(\emptyset))+(v(\pi(1)\pi(2))-v(\pi(1)))$
$+(v(\pi(1)\pi(2)\pi(3))-v(\pi(1)\pi(2)))+$
$\ldots+(v(\pi(1)\pi(2)\pi(3)\cdots\pi(|P(\pi)|))$
$-v(\pi(1)\pi(2)\cdots\pi(|P(\pi)|-1)))=v(\pi).$
Where $\pi(k)$ is the agent at the $k^{\textup{th}}$ position of $\pi$.
For each carrier game, we know the structure of the unique solution concept that satisfies SE and SNP from Lemma˜4. Note that MargSol satisfies all three properties for every game $\langle N,v\rangle$, hence it must satisfy them for the carrier games as well. Suppose there exists a different solution concept $\phi$ that also satisfies the same three properties. We will prove that $\phi$ must be MargSol.
From Lemma˜4 and the discussion above, we conclude that $\phi$ must be the same as MargSol for the carrier game $\langle N,u_{\pi,\alpha}\rangle$. Hence we have $\forall i\in N,\forall\pi,\pi^{\prime}\in\Pi$
$\textsc{MargSol}_{i}(\pi^{\prime},u_{\pi,\alpha_{\pi}})=\phi_{i}(\pi^{\prime},u_{\pi,\alpha_{\pi}}),$
$\Rightarrow\sum_{\pi\in\Pi}\textsc{MargSol}_{i}(\pi^{\prime},u_{\pi,\alpha_{\pi}})=\sum_{\pi\in\Pi}\phi_{i}(\pi^{\prime},u_{\pi,\alpha_{\pi}}),$
$\Rightarrow\textsc{MargSol}_{i}\left(\pi^{\prime},\sum_{\pi\in\Pi}u_{\pi,\alpha_{\pi}}\right)=\phi_{i}\left(\pi^{\prime},\sum_{\pi\in\Pi}u_{\pi,\alpha_{\pi}}\right),$
$\Rightarrow\textsc{MargSol}_{i}\left(\pi^{\prime},v\right)=\phi_{i}\left(\pi^{\prime},v\right).$
The first implication holds by summing over all $\pi\in\Pi$. The second implication is by SA for both solution concepts. The third implication is by Equation˜5. Since $v$ was arbitrary, we conclude that $\phi$ is the same as MargSol. ∎
B.4 Proof of Theorem˜3
To prove this result, we first need the following lemma.
Lemma 6.
Proof.
Now we present the proof of Theorem˜3.
Proof.
($\Rightarrow$) This direction of the proof follows in similar lines to that of Theorem˜2. Assume that there exists an extended solution concept $\psi\neq\textsc{Ext-Shap}{}$ that also satisfies the three given properties.
For each carrier game, we know the structure of the unique solution concept that satisfies SE and SNP from Lemma˜6. Note that Ext-Shap satisfies all three properties for every game $\langle N,v\rangle$, hence it must satisfy them for the carrier games as well. Hence $\psi$ must be same as Ext-Shap for the carrier game $\langle N,u_{\pi,\alpha}\rangle$. Hence we have
The first implication holds by summing over all $\pi\in\Pi$. The second implication is by EA for both solution concepts. The third implication is by Equation˜5. Since $v$ was arbitrary, we conclude that $\psi$ is the same as Ext-Shap. ∎
Appendix C Appendix to Section˜7
C.1 Proof of Lemma˜5
Proof.
$\phi_{i}(\pi)\leqslant\phi_{i}(\pi^{*}(v)),\ \forall\pi\in\Pi$
Averaging over all possible $\pi\in\Pi$ such that $P(\pi)=N$, we get
$\overline{\phi}_{i}=$ $\frac{\textstyle 1}{\textstyle n!}\sum_{\pi\in\Pi:P(\pi)=N}\phi_{i}(\pi)$ (6)
$<$ $\frac{\textstyle 1}{\textstyle n!}\sum_{\pi\in\Pi:P(\pi)=N}v(\pi_{i}+i)-v(\pi_{i})=\textsc{Ext-Shap}_{i}.$ (7)